OpenStudy (anonymous):

find the percent of the change in these numbers 20 to 17

ganeshie8 (ganeshie8):

use this : % change = (new - old)/old * 100 = (17-20)/20 * 100 = .15 * 100 = 15 %

OpenStudy (anonymous):

incease or decease

ganeshie8 (ganeshie8):

what do you think. going from 20 to 17 is increase or decrease ?

OpenStudy (anonymous):

ohh decrease

ganeshie8 (ganeshie8):

yes !